# dc-GAN: Dual-Conditioned GAN for Face Demorphing From a Single Morph

### v2

@@ 5.1 Evaluation Criteria
For differential demorphing, we use a modified approach. Given a MORPH and a reference image REF, our method generates an output image, OUT. We treat the ground-truth image BF and OUT as the genuine pair. For the imposter pair, we search the entire test set of bonafide images in the FRLL dataset and assign the face with the third highest similarity to OUT, excluding GT and REF for obvious reasons (hence third). During evaluation, we arbitrarily select one of the constituent face image used to create MORPH as the reference image. However, this does not reflect a real-world scenario, as the exact reference image used to create the morph might not be available during demorphing. To simulate this, we apply one or more random transformations to the reference image (see Figure 3), such as Gaussian blur, jitter, affine transformation, etc. It is important to note that these transformations were not applied during training.
Table 3: Reference Free Demorphing: We compute the True Match Rate (TMR in %) at various False Match Rate (FMR in %) thresholds (0.1/1/5/10) between the generated outputs (OUT1, OUT2) and the ground truth images (BF1, BF2) across the five FR models used. ArcFace performs most consistently across the FRS used in the presence of high-frequency artefacts (See Figure 4 and 5)
@@
Table 4: Differential Demorphing: We compute the True Match Rate (TMR in %) at various False Match Rate (FMR in %) thresholds between the generated output (OUT) and the ground truth image (GT) across the five FR models used.
@@
Table 5: Comparison of our reference-free approach to SDeMorph [38] and Identity-Preserving Demorphing (IPD) [39]. Unlike our method, which operates on unseen faces (scenario 3) and is significantly more challenging, the other methods assume the same identities are present in both training and test morphs (scenario 1). In scenario 3, our method significantly outperforms the current state-of-the-art method. The scores for scenario 3 are provided by the original authors of [38, 39].
@@
Figure 4: Reference-Free Demorphing Visualization: We illustrate the results of our reference-free demorphing approach. The model takes the morph image (MORPH) as input and produces two outputs (OUT1 and OUT2). The ground truth images used to generate the morph are denoted as BF1 and BF2.
Figure 5: Differential Demorphing Visualization: We illustrate the results of our differential demorphing approach. The model receives the morph image (MORPH) and the transformed ground-truth image (TRANSFORMED BF1) and produces an output (OUT). The ground truth (GT) is shown, which, along with the ground-truth image (BF1), was used to generate the morph image (MORPH). Note that our method does not assume the demorphing process uses the exact same ground-truth image (BF1) that was used to create the morph.
Table 6: We compute image quality between generated face images and ground truth face images.
@@ 5.2 Findings
We evaluate our method both visually and analytically. It is imperative for any demorphing method to produce visually distinctive faces compared to the input morph as well as between themselves. Note that a trivial solution for demorphing would be an identity function which simply outputs the morph. By construction, the outputs will have the best matches with the bonafide faces used to create the morph in the database, rendering the demorphing process unnecessary. Thus, visual evaluation is paramount in demonstrating the effectiveness of our method.
5.2.1 Reference Free Demorphing
@@
We visualize the outputs of our reference-free demorphing method across six morphing techniques in Figure 4. These techniques include both landmark-based and deep-learning-based morphs. The first three columns visualize the morph and the two constituent face images used to create it, while the subsequent two columns show the outputs of our method, presented in no specific order. We observe that the faces produced by our method are visually distinct from each other and also appear notably different from the original morph. Moreover, they are visually closer to their ground-truth counterparts (refer to Appendix for more images). To quantify this, we calculate the normalized biometric distance between the generated faces and between the generated faces and the ground truth. We use five well-known face biometric models and plot the distance scores across three different pairings of images: OUT1-OUT2, BF1-OUT1, and BF2-OUT2, which represent the dissimilarity between the generated outputs and between each generated output and its corresponding ground truth. We plot the distance distribution for all datasets across all FR models in Figure 6. We observe an overall OUT1-OUT2 biometric distance of 0.4380, indicating a quantifiable difference in the output faces. Note that we use cosine distance rather than similarity to quantify results. This choice helps us intuitively explain morph-replication because a greater distance indicates that the generated faces are more distinguishable from each other. For BF1-OUT1, we observe a distance of 0.3992, and for BF2-OUT2, a distance of 0.4025. Note that the distance scores between outputs and their respective ground-truth face images fall well below the threshold used in several biometric systems [37, 36]. Our method successfully learns to distinguish features from the base image, such as hair (Rows 1 and 3), skin color, and subtle details like eyebrows and facial hair. Finally, we compute the True Match Rate (TMR) of the output subjects at various False Match Rate (FMR) thresholds and present the results in Table 3. Our method achieves an average TMR of 93.86% on the AMSL dataset. The corresponding TMR values on the OpenCV, StyleGAN, WebMorph, FaceMorph, and MorDiff datasets are 93.99%, 58.92%, 89.87%, 94.39%, and 93.75%, respectively. We also evaluate the quality of images generated by our method using various image-quality assessment (IQA) metrics, as shown in Table 6. Our method achieves an average FID [17] of 0.4284 across all datasets, with corresponding averages of 0.61 for SSIM [41], 0.24 for LPIPS [43], and 18.02 for PSNR, demonstrating the visual realism of the outputs from our method. We also observe that ArcFace performs consistently across all datasets in the presence of visual artefacts generated during demorphing compared to other FRS which is consistent with [5].
@@
Comparison with SOTA: We evaluate our method against existing reference-free demorphing approaches based on restoration accuracy and TMR at 10% FMR. Restoration accuracy measures the proportion of outputs that correctly match their corresponding ground-truth images. Our method achieves an average TMR of 93.86% on the AMSL dataset using the ArcFace FRS, surpassing the performance of [5] by 23.32%. Reference-free demorphing methods, particularly for scenario 3, remain limited. Nonetheless, we also compare our method’s restoration accuracy with [38, 39] in both scenarios 1 and 3. In scenario 3, our method shows a significant performance advantage and achieves comparable results in scenario 1, despite being trained in the more challenging scenario 3. Results are summarized in Table 5.
@@
Effect of $\mathcal{L}_{cr}$ and $\mathcal{E}(\cdot)$: We conduct an ablation study on our method to assess i) the impact of cross-road loss and ii) the influence of dual conditioning on generation. Excluding pixel-wise guidance (the second part of the loss in Equation 9) results in an average performance decline of 20.87% across the six datasets tested on the ArcFace FRS. This decrease is mainly due to the model’s reduced ability to accurately align the two unordered generated outputs with the ordered ground-truth images. Moreover, when the generator is trained without $\mathcal{E}$, the performance drops by 8.57% due to lack of context during the generation (see Section 3). These results are presented in Table 9.
5.2.2 Differential demorphing
@@
We visualize the results of differential demorphing in Figure 5. The first two columns show MORPH and BF1, the third column displays the randomly transformed BF1 used as the reference image in the demorphing process, and the remaining two columns present the output (OUT) from our method and the ground truth (GT). Our method produces faces that are visibly distinct from both the MORPH and the reference image achieving a TMR of 94.88% on the AMSL dataset, with corresponding values of 88.12%, 54.04%, 89.38%, 98.0%, and 90.32% on the OpenCV, StyleGAN, WebMorph, FaceMorph, and MorDiff datasets, respectively. These results are presented in Table 4.
@@
Table 7: We calculate the average biometric distance for three cases —OUT1-OUT2, BF1-OUT1, and BF2-OUT2—across all five FR models utilized. (Individual scores in Appendix.)
@@
Table 8: We calculate the $d^{\prime}$ values for True versus Imposter scores in differential demorphing. Higher $d^{\prime}$ values suggest greater separability between the distributions.
@@
Table 9: Effect of the dual-condition and $\mathcal{L}_{cr}$ across five FRS used (AdaFace/ArcFace/VGG-Face/FaceNet/OpenFace). The drop in TMR performance shows the importance of the pixel-wise guidance during training and the effectiveness of the condition by $\mathcal{E}$ during generation.
6 Summary
In this paper, we propose dc-GAN, a novel method to recover facial identities from a morph. Our proposed method is based on GAN architecture, we condition our GAN on morphs as well as their latent representation to provide robust guidance while recovering identities. Previous approaches suffered from morph-replication, i.e., they produced similar-looking faces. Our method overcomes this problem significantly. We evaluate our method on 6 different morphing techniques using five biometric face models. We assessed our method using the AMSL, FRLL-Morphs, and MorDiff datasets, achieving visually compelling results and addressing the issue of morph-replication. Moreover, experiments suggest that ArcFace performs most consistently for demorphing tasks. Future work will aim to develop techniques for demorphing faces in various styles, beyond just passport-style face images.

### v4

@@ V-A Evaluation Criteria
For differential demorphing, we use a modified approach. Given a MORPH and a reference image REF, our method generates an output image, OUT. We treat the ground-truth image BF and OUT as the genuine pair. For the imposter pair, we search the entire test set of bonafide images in the FRLL dataset and assign the face image with the third highest similarity to OUT, excluding GT and REF for obvious reasons (hence the third). During evaluation, we arbitrarily select one of the constituent face images used to create MORPH as the reference image. However, this does not reflect a real-world scenario, as the exact reference image used to create the morph might not be available during demorphing. To simulate this, we apply one or more random transformations to the reference image (see Figure 3), such as Gaussian blur, jitter, affine transformation, etc. It is important to note that these transformations were not applied during training.
TABLE III: Reference-Free Demorphing: We compute the True Match Rate (TMR in %) at various False Match Rate (FMR in %) thresholds (0.1/1/5/10) between the generated outputs (OUT1, OUT2) and the ground truth images (BF1, BF2) across the five FR models used. ArcFace performs most consistently across the various models used in the presence of high-frequency artifacts (See Figure 4 and 5)
@@
TABLE IV: Differential Demorphing: We compute the True Match Rate (TMR in %) at various False Match Rate (FMR in %) thresholds between the generated output (OUT) and the ground-truth image (GT) across the five FR models used.
@@
TABLE V: Restoration Accuracy: Comparison of our reference-free approach to SDeMorph [38] and Identity-Preserving Demorphing (IPD) [39]. Unlike our method, which operates on unseen faces (scenario 3) and is significantly more challenging, the other methods assume the same identities are present in both training and test morphs (scenario 1). In scenario 3, our method significantly outperforms the current state-of-the-art method. The scores for scenario 3 are provided by the original authors of [38, 39].
@@
Figure 4: Reference-Free Demorphing Visualization: We illustrate the results of our reference-free demorphing approach. The model takes the morph image (MORPH) as input and produces two outputs (OUT1 and OUT2). The ground-truth images used to generate the morph are denoted as BF1 and BF2.
Figure 5: Differential Demorphing Visualization: We illustrate the results of our differential demorphing approach. The model receives the morph image (MORPH) and the transformed ground-truth image (TRANSFORMED BF1) and produces an output (OUT). The ground truth (GT) is shown, which, along with the ground-truth image (BF1), was used to generate the morph image (MORPH). Note that our method does not assume that the demorphing process uses the exact same ground-truth image (BF1) that was used to create the morph.
TABLE VI: We compute image quality between generated face images and ground truth face images.
@@ V-B Findings
We evaluate our method both visually and analytically. It is imperative for any demorphing method to produce visually distinctive faces compared to the input morph as well as between themselves. Note that a trivial solution for demorphing would be an identity function which simply outputs the morph. By construction, the outputs will have high similarity scores with the bonafide faces used to create the morph in the database, rendering the demorphing process unnecessary. Therefore, visual evaluation plays a crucial role in demonstrating the effectiveness of our method.
V-B1 Reference-Free Demorphing
@@
We visualize the outputs of our reference-free demorphing method across six morphing techniques in Figure 4. These techniques include both landmark-based and deep-learning-based morphs. The first three columns visualize the morph and the two constituent face images used to create it, while the subsequent two columns show the outputs of our method, presented in no specific order. We observe that the faces produced by our method are visually distinct from each other and also appear notably different from the original morph. Moreover, they are visually closer to their ground-truth counterparts. To quantify this, we calculate the normalized biometric distance between the generated faces and between the generated faces and the ground truth. We plot the distance scores across three different pairings of images: OUT1-OUT2, BF1-OUT1, and BF2-OUT2, which represent the dissimilarity between the generated outputs, and between each generated output and its corresponding ground-truth. We plot the distance distribution for all datasets across all FR models in Figure 6. We observe an overall OUT1-OUT2 biometric distance of 0.4380, indicating a quantifiable difference in the output faces. Note that we use cosine distance rather than similarity to quantify results. This choice helps us intuitively explain morph-replication because a greater distance indicates that the generated faces are more distinguishable from each other. For BF1-OUT1, we observe a distance of 0.3992, and for BF2-OUT2, a distance of 0.4025. Note that the distance scores between outputs and their respective ground-truth face images fall well below the threshold used in several biometric systems [37, 36]. Our method successfully learns to distinguish features from the base image, such as hair (Rows 1 and 3), skin color, and subtle details like eyebrows and facial hair. Finally, we compute the True Match Rate (TMR) of the output subjects at various False Match Rate (FMR) thresholds and present the results in Table III. Our method achieves an average TMR of 93.86% on the AMSL dataset. The corresponding TMR values on the OpenCV, StyleGAN, WebMorph, FaceMorph, and MorDiff datasets are 93.99%, 58.92%, 89.87%, 94.39%, and 93.75%, respectively. We also evaluate the quality of images generated by our method using various image quality assessment (IQA) metrics, as shown in Table VI. Our method achieves an average FID [17] of 0.4284 across all datasets, with corresponding averages of 0.61 for SSIM [41], 0.24 for LPIPS [43], and 18.02 for PSNR, demonstrating the visual realism of the outputs produced by our method. We also observe that ArcFace performs consistently across all datasets - despite the presence of visual artifacts generated during demorphing - compared to other face models, which is in line with [5].
@@
Comparison with SOTA: We evaluate our method against existing reference-free demorphing approaches based on restoration accuracy and TMR at 10% FMR. Restoration accuracy measures the proportion of outputs that correctly match their corresponding ground-truth images. Our method achieves an average TMR of 93.86% on the AMSL dataset using the ArcFace model, surpassing the performance of [5] by 23.32%. Reference-free demorphing methods, particularly for scenario 3, remain limited. Nonetheless, we compare our method’s restoration accuracy with [38, 39] in both scenarios 1 and 3. In scenario 3, our method shows a significant performance advantage and achieves comparable results in scenario 1, despite being trained in the more challenging scenario 3. Results are summarized in Table V.
@@
Effect of $\mathcal{L}_{cr}$ and $\mathcal{E}(\cdot)$: We conduct an ablation study on our method to assess i) the impact of cross-road loss and ii) the influence of dual conditioning on output generation. Excluding pixel-wise guidance (the second part of the loss in Equation 8) results in an average performance decline of 20.87% across the six datasets tested using the ArcFace model. This decrease is mainly due to the method’s reduced ability to accurately align the two unordered generated outputs with the ordered ground-truth images. Moreover, when the generator is trained without $\mathcal{E}$, the performance drops by 8.57% due to lack of context during the generation process (see Section III). These results are presented in Table IX.
V-B2 Differential demorphing
@@
We visualize the results of differential demorphing in Figure 5. The first two columns show MORPH and BF1, the third column displays the randomly transformed BF1 used as the reference image in the demorphing process, and the remaining two columns present the output (OUT) from our method and the corresponding ground truth (GT). Our method produces faces that are visibly distinct from both the MORPH and the reference image, achieving a TMR of 94.88% on the AMSL dataset, with corresponding values of 88.12%, 54.04%, 89.38%, 98.0%, and 90.32% on the OpenCV, StyleGAN, WebMorph, FaceMorph, and MorDiff datasets, respectively. These results are presented in Table IV.
@@
TABLE VII: We calculate the average biometric distance for three cases — OUT1-OUT2, BF1-OUT1, and BF2-OUT2 — across all five FR models utilized.
@@
TABLE VIII: We calculate the $d^{\prime}$ values for Genuine versus Imposter scores in differential demorphing. Higher $d^{\prime}$ values suggest greater separability between the distributions.
@@
TABLE IX: Effect of the dual-condition and $\mathcal{L}_{cr}$ across the five FR models used (AdaFace/ArcFace/VGG-Face/FaceNet/OpenFace). The drop in TMR performance shows the importance of the pixel-wise guidance during training and the effectiveness of the condition by $\mathcal{E}$ during output generation.
@@
VI Summary
In this paper, we propose dc-GAN, a novel method to recover constituent images from a facial morph. Our proposed method is based on the GAN architecture; we condition our GAN on morphs as well as their latent representations to provide robust guidance while recovering constituent images. Previous approaches suffered from morph-replication, i.e., they produced similar looking faces as the morph itself. Our method significantly overcomes this problem. We evaluate our method across 6 different morphing techniques using five biometric face models. We assess our method using the AMSL, FRLL-Morphs, and MorDiff datasets, achieving visually compelling results and addressing the issue of morph-replication. Moreover, experiments suggest that ArcFace performs consistently for demorphing tasks compared to other face recognition models. Future work will aim to develop techniques for demorphing faces with various styles, beyond just passport-style face images.
VII Acknowledgment
This work was supported in part by the NSF Center for Identification Technology Research (CITeR) and the DHS Center for Criminal Investigations and Network Analysis (CINA).
This research did not involve the collection of data from human subjects. All experiments were carried out using existing face datasets available for research purposes, and no attempts were made to identify individuals within these datasets. Additionally, no personally identifiable information (PII) was used during the demorphing process described in this paper.